### v25

# The Riemann Hypotheses is false

ABSTRACT. Let $\Theta$ denote the supremum of the real parts of the zeros of the Riemann zeta function. We demonstrate that $\Theta=1$, which entails the existence of infinitely many Riemann zeros off the critical line (thus disproving the Riemann Hypothesis (RH)). The paper is concluded by a brief discussion of why our argument doesn’t work for both Weil and Beurling zeta functions whose analogues of the RH are known to be true.
2020 Mathematics Subject Classifications: 11M26, 11M06.
Keywords and phrases: Riemann zeta function, Riemann Hypothesis, disproof.
Introduction. The Riemann zeta function is a function of the complex variable $s$, defined in the half-plane $\Re(s)>1$ by $\zeta(s):=\sum_{n=1}^{\infty}n^{-s}$ and in the whole complex plane by holomorphic continuation. Euler noticed that for $\Re(s)>1$, $\zeta(s)$ can be expressed as a product $\prod_{p}(1-p^{-s})^{-1}$ over the entire set of primes, which entails that $\zeta(s)\neq 0$ for $\Re(s)>1$. It can be shown that $\zeta(s)$ extends to $\mathbb{C}$ as a meromorphic function with only a simple pole at $s=1$, with residue $1$, and satisfies the functional equation $\xi(s)=\xi(1-s)$, where $\xi(z)=\frac{1}{2}z(z-1)\pi^{-z/2}\Gamma(\frac{1}{2}z)\zeta(z)$ and $\Gamma(w)=\int_{0}^{\infty}e^{-x}x^{w-1}\mathrm{d}x$. Let $\chi$ be a primitive Dirichlet character and $L(s,\chi)$ be the corresponding L-function. Most of the analytic properties for $\zeta$ are also known to hold for $L(s,\chi)$ [5, Chapter 10]. Let $\Theta_{\chi}$ be the supremum of the real parts of the zeros of $L(s,\chi)$. The Generalised Riemann Hypothesis (GRH) is equivalent to the statement that $\Theta_{\chi}=\frac{1}{2}$ for each $\chi$, and the Generalised Prime Number Theorem is equivalent to the statement that $L(s,\chi)\neq 0$ at $\Re(s)=1$ for each $\chi$. For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
Definitions. Let: $\rho$ denote a non-trivial zero of the Riemann zeta function, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s. Denote by $\Lambda$ the von Mangoldt function.
Theorem 1. One has $\Theta=1$.
Proof.
Suppose that $\Theta<1$ and fix $0<\varepsilon<1-\Theta$. Let $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$. By Lemma 12.2 of [5], we have $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX)\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$. Thus we can choose $T$ and $H\in[3T,3T+1]$ such that
$\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+2iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
uniformly for $-1\leq\sigma\leq 2$. By the residue theorem, we have
$\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{\Theta+\varepsilon+2iT}^{\Theta+\varepsilon+iH}+\int_{\Theta+\varepsilon+iH}^{-1+iH}+\int_{-1+iH}^{-1+2iT}+\int_{-1+2iT}^{\Theta+\varepsilon+2iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
$=I_{1}+I_{2}+I_{3}+I_{4},$ (3)
say. From now on, let $2T\leq t\leq H$ and $k=k(\varepsilon)>4$ be a fixed real number such that $k(1-\Theta-\varepsilon)>2$. For $0<a<b\leq 2a$, note that
$\sum_{a\leq n\leq b}n^{it}\ll t^{1/2}+at^{-1/2}$ (4)
[6, p.107]. Hence
$\sum_{\frac{T}{\log^{k}T}<n\leq T}n^{it}$ $=\Big{(}\sum_{\frac{T}{\log^{k}T}<n\leq\frac{2T}{\log^{k}T}}+\sum_{\frac{2T}{\log^{k}T}<n\leq\frac{4T}{\log^{k}T}}+\cdots+\sum_{\frac{T}{2}<n\leq T}\Big{)}n^{it}$ (5)
$\ll T^{\frac{1}{2}}\log^{k+1}T.$ (6)
By partial summation, one deduces from (6) that
$\sum_{\frac{T}{\log^{k}T}<n\leq T}n^{\sigma-1+it}\ll T^{\sigma-\frac{1}{2}}\log^{k+1}T$ (7)
for $\sigma\in\mathbb{R},|\sigma|\leq A$ where $A>0$ is a constant. Combining (1) and (7) yields $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}.$ Let $\sigma\geq\varepsilon$. By the log-derivative of the functional equation for $\zeta$ (see e.g. [3, p.403]), note that
$\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log t$ (8)
uniformly for $t\geq 2$. Combining (7) and (8) gives $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{k+2}T.$ Thus we arrive at
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{k+2}T.$ (9)
Let $s=\Theta+\varepsilon+iY$ and $c=1-\Theta-\varepsilon+\frac{1}{\log T}>0$, where $Y\in[2T,H]$. By Theorem 3.12 of [6], note that
$-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}=\frac{1}{2\pi i}\int_{c-iT}^{c+iT}\frac{\zeta^{\prime}}{\zeta}(s+w)\frac{(T/2\log^{k}T)^{w}}{w}\mathrm{d}w+O(T^{c-1}\log^{1-kc}T).$ (10)
Notice that the above integral has a simple pole at $w=1-s$ with residue $\frac{\zeta^{\prime}}{\zeta}(s)-\frac{(T/2\log^{k}T)^{1-s}}{1-s}.$ Hence, shifting the line of integration in (10) to $\Re(s)=-1$ then invoking (1) and (8) to bound the integrals from Cauchy’s residue theorem, gives
$\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\Lambda(m)}{m^{s}}-\frac{(T/2\log^{k}T)^{1-s}}{1-s}+O(T^{c-1}\log^{1-kc}T).$ (11)
Putting (11) into $I_{1}$ gives
$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\sum_{m\leq\frac{T}{2\log^{k}T}}\Bigg{(}\frac{n^{\Theta+\varepsilon-1+iT}\Lambda(m)}{m^{\Theta+\varepsilon+iT}\log(n/m)}-\frac{n^{\Theta+\varepsilon-1+iH}\Lambda(m)}{m^{\Theta+\varepsilon+iH}\log(n/m)}\Bigg{)}$ (12)
$-\frac{1}{2\pi i}\int_{\Theta+\varepsilon+iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s+O(n^{\Theta+\varepsilon-1}T^{c}\log^{2-kc}T).$ (13)
Since $1\leq m\leq\frac{T}{2\log^{k}T}$, note that $0<\frac{1}{\log(n/m)}\leq\frac{1}{\log 2}$ uniformly for $\frac{T}{\log^{k}T}\leq n\leq T$. Hence by partial summation, it follows from (6) that
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iT}}{m^{\Theta+\varepsilon+iT}\log(n/m)}\ll\frac{T^{\frac{1}{2}}\log^{k+1}T}{m^{\Theta+\varepsilon}\log(T/m)}$ (14)
uniformly for $2\leq m\leq\frac{T}{2\log^{k}T}$. The same bound holds for $\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{n^{iH}}{m^{\Theta+\varepsilon+iH}\log(n/m)}.$ By partial summation, it also follows from (6) that
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\int_{\Theta+\varepsilon+iT}^{\Theta+\varepsilon+iH}\frac{(T/2n\log^{k}T)^{1-s}}{1-s}\mathrm{d}s\ll T^{1/2}\log^{k+1}T.$ (15)
Recall that $H=3T+O(1)$ and $kc>2$. Thus, summing (12) and (13) for $\frac{T}{\log^{k}T}\leq n\leq T$ yields
$\sum_{\frac{T}{\log^{k}T}\leq n\leq T}I_{1}$ $=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{\Lambda(n)}{n}+O\Bigg{(}T^{\frac{1}{2}}\log^{k+1}T\sum_{m\leq\frac{T}{2\log^{k}T}}\frac{\log m}{m^{\Theta+\varepsilon}\log(T/m)}\Bigg{)}+o(T)$ (16)
$=-\frac{T}{2\pi}\sum_{\frac{T}{\log^{k}T}\leq n\leq T}\frac{\Lambda(n)}{n}+o(T)$ (17)
$=-\frac{kT}{2\pi}\log\log T+o(T)$ (18)
since $\sum_{m<x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ [5, p.182], where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant. By Theorem 1 of [3], we have
$\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (19)
uniformly for $2\leq n\ll T.$ Define $N(y):=\sum_{0<\gamma\leq y}1$. Note that $N(y)=(y/2\pi)\log(y/2\pi)-(y/2\pi)+O(\log y)$ [5, Corollary 14.3]. Thus from (19) we obtain
$\sum_{1\leq n\leq\frac{T}{\log^{k}T}}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(2T)-\frac{T}{2\pi}\sum_{2\leq n\leq\frac{T}{\log^{k}T}}\frac{\Lambda(n)}{n}+o(T)$ (20)
$=\frac{kT}{2\pi}\log\log T+\frac{T}{2\pi}(B-1+\log 2)+o(T),$ (21)
since $H=3T+O(1)$. Combining (21), (18) and (9) gives
$\sum_{1\leq n\leq T}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\sum_{2T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}$ (22)
$=\Big{(}\sum_{2T\leq\gamma_{T}\leq H}\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{k}T}}+\sum_{\frac{T}{\log^{k}T}<n\leq T}\Big{)}\Big{)}n^{\rho_{T}-1}$ (23)
$\gg T.$ (24)
By Theorem 4.11 of [6], we have
$\sum_{n\leq T}n^{\rho_{T}-1}=\zeta(1-\rho_{T})+\frac{T^{\rho_{T}}}{\rho_{T}}+O(T^{\Theta-1})\ll T^{\Theta-1}$ (25)
uniformly for $2T\leq\gamma_{T}\leq H$. Hence
$\sum_{1\leq n\leq T}\sum_{2T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{2T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (26)
But (26) is a contradiction to (24), thus we deduce that our supposition must be false. ∎
Corollary 1. There exists an infinite sequence of zeros of $\zeta$, whose real parts converge to 1.
Proof.
Combine Theorem 1 with the fact that $\zeta(s)\neq 0$ for $\Re(s)\geq 1$ [5, Theorem 6.6]. ∎
Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $12T\leq\gamma_{T}\leq 3T+$ 1 and $\Re(\rho_{T})=1-o(1)$.
Remark 2. Our argument crucially relies on:
(i) The functional equation for $\zeta$,
(ii) Theorem 3.12 of [6], and
(iii) Theorem 4.11 of [6].
Both (i) and (iii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers. Hence our approach wouldn’t work for such, whose analogues of the RH are known to be true. Note that both (i) and (ii) also do not hold for the Weil zeta functions, which are also known to satisfy the RH.