### v38

# The Riemann Hypothesis is false

ABSTRACT. Let $\Theta$ denote the supremum of the real parts of the zeros of the Riemann zeta function. We demonstrate that $\Theta=1$, which entails the existence of infinitely many Riemann zeros off the critical line (thus disproving the Riemann Hypothesis (RH)). The paper is concluded by a brief discussion of why our argument doesn’t work for both Weil and Beurling zeta functions whose analogues of the RH are known to be true.
2020 Mathematics Subject Classifications: 11M26, 11M06.
Keywords and phrases: Riemann zeta function, Riemann Hypothesis, disproof.
 [^1] Data availability statement: This manuscript has no associated data. Introduction. The Riemann zeta function is a function of the complex variable $s$, defined in the half-plane $\Re(s)>1$ by $\zeta(s):=\sum_{n=1}^{\infty}n^{-s}$ and in the whole complex plane by holomorphic continuation. Euler noticed that for $\Re(s)>1$, $\zeta(s)$ can be expressed as a product $\prod_{p}(1-p^{-s})^{-1}$ over the entire set of primes, which entails that $\zeta(s)\neq 0$ for $\Re(s)>1$. It can be shown that $\zeta(s)$ extends to $\mathbb{C}$ as a meromorphic function with only a simple pole at $s=1$, with residue $1$, and satisfies the functional equation $\xi(s)=\xi(1-s)$, where $\xi(z)=\frac{1}{2}z(z-1)\pi^{-z/2}\Gamma(\frac{1}{2}z)\zeta(z)$ and $\Gamma(w)=\int_{0}^{\infty}e^{-x}x^{w-1}\mathrm{d}x$. Let $\chi$ be a primitive Dirichlet character and $L(s,\chi)$ be the corresponding L-function. Most of the analytic properties for $\zeta$ are also known to hold for $L(s,\chi)$ [5, Chapter 10]. Let $\Theta_{\chi}$ be the supremum of the real parts of the zeros of $L(s,\chi)$. The Generalised Riemann Hypothesis (GRH) is equivalent to the statement that $\Theta_{\chi}=\frac{1}{2}$ for each $\chi$, and the Generalised Prime Number Theorem is equivalent to the statement that $L(s,\chi)\neq 0$ at $\Re(s)=1$ for each $\chi$. For a far more thorough discussion of the RH and GRH, the interested reader is kindly referred to Borwein et al [1].
Definitions. Let: $\rho$ denote a non-trivial zero of $\zeta$, $\gamma=\Im(\rho)$ and $\Theta\in[\frac{1}{2},1]$ be the supremum of the real parts of the $\rho$’s.
Lemma 1. (Theorem 4.11 of [7]). We have $\zeta(s)=\sum_{n\leq x}n^{-s}-\frac{x^{1-s}}{1-s}+O(x^{-\sigma})$ uniformly for $\sigma\geq\sigma_{0}>0,|\Im(s)|<2\pi x/C$, when $C$ is a given constant $>1$.
Lemma 2, (Theorem 1 of [3]). One has $\sum_{T\leq\gamma_{T}\leq 2T+1}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ uniformly for $2\leq n\ll T.$
Theorem 1. One has $\Theta=1$.
Proof.
Let: $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$. By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$. Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
$\max\Big{(}\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iT)\Big{|},\Big{|}\frac{\zeta^{\prime}}{\zeta}(\sigma+iH)\Big{|}\Big{)}\ll\log^{2}T$ (1)
uniformly for $-1\leq\sigma\leq 2$. By the residue theorem, we have
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+\varepsilon+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+\varepsilon+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{\Theta+\varepsilon+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
$=I_{1}+I_{2}+I_{3}+I_{4},$ (3)
say. From now on, let $T\leq t\leq H$ and $\sigma=\Re(s)$. By the approximate functional equation for $\zeta$ (Theorem 4.15 of [7]), note that
$\zeta(\sigma-it)=\sum_{n\leq x}n^{-\sigma+it}+\chi(\sigma-it)\sum_{v\leq y}v^{\sigma-1-it}+O(x^{-\sigma}+t^{\frac{1}{2}-\sigma}y^{\sigma-1})$ (4)
where $00\leq\sigma\leq 1,x>h>0,y>h>0,t=2\pi xy,h>$ 0 is a constant and $|\chi(\sigma-it)|=\Big{|}\Big{(}\frac{t}{2\pi}\Big{)}^{\sigma-\frac{1}{2}-it}e^{i(t+\frac{\pi}{4})}(1+O(1/t))\Big{|}\sim(t/2\pi)^{\frac{1}{2}-\sigma}$ [7, p.78]. Thus since $|\zeta(-it)|\asymp t^{1/2}|\zeta(1-it)|\ll T^{1/2}\log T$, it follows by taking $\sigma=0$ in (4) that
$S(x):=\sum_{n\leq x}n^{it}\ll T^{\frac{1}{2}}\log^{3}T$ (5)
uniformly for $\frac{T}{\log^{2}T}\leq x\leq T$. Let $f(n)=n^{\sigma-1}.$ Then by Abel summation, it follows that
$\sum_{n\leq x}n^{\sigma-1+it}=S(T)f(T)-S(T/\log^{2}T)f(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)f^{\prime}(x)\mathrm{d}x\ll T^{\sigma-\frac{1}{2}}\log^{7}T$ (6)
for $-1\leq\sigma\leq 2.$ Combining (1) and (6) gives $\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{7}T.$ By the log-derivative of the functional equation for $\zeta$, we have
$\frac{\zeta^{\prime}}{\zeta}(-\sigma+it)\ll\log T$ (7)
for $\sigma\geq\sigma_{0}>0$. Combining (6) and (7) yields $\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{7}T.$ Thus we arrive at
$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{7}T.$ (8)
Note that $\frac{\zeta^{\prime}}{\zeta}(s)=-\sum_{m=1}^{\infty}\frac{\Lambda(m)}{m^{s}}$ for $\Re(s)=\sigma>1$. Inserting this into $I_{1}$ gives
$I_{1}=\frac{(T-H))\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=1,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{\log(n/m)}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. Let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=-\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}$ $=S(T)g(T)-S(T/\log^{2}T)g(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
$\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}-T^{1/2}\log^{3}T\int_{T/\log^{2}T}^{T}g^{\prime}(x)\mathrm{d}x$ (11)
$\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (12)
uniformly for $1\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon).$ For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll 1/\varepsilon=T^{1/4}$ so it similarly follows from (5) by Abel summation that
$\sum_{\frac{T}{\log^{2}T}\leq n\leq\frac{m}{1+\varepsilon}\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (13)
and
$\sum_{\frac{T}{\log^{2}T}\leq m(1+\varepsilon)\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (14)
We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$. For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=r/m(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=m/r(1+O(r/m))$ hence
$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}$ $=\frac{\Lambda(m)}{m^{it}}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}$ (15)
$=\frac{\Lambda(m)}{em^{it}}\sum_{0<r\leq\varepsilon m-1}\Big{(}\frac{(m+r)^{it}}{r}+O(r/m^{2})\Big{)}$ (16)
$=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq\varepsilon m-1}\frac{(1+\frac{r}{m})^{it}}{r}+O(\varepsilon^{2}\log m).$ (17)
Note that if $f(t)$ is differentiable on $[a,b]$, $f^{\prime}(t)$ is monotonic decreasing on $[a,b]$, $f^{\prime}(a)=\beta,f^{\prime}(b)=\alpha,$ then
$\sum_{a<r\leq b}e^{2\pi if(r)}=\sum_{|v|\leq N}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x+o(1)$ (18)
as $N\rightarrow\infty$, where
$\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x\ll\Big{|}\frac{1}{\alpha-v}-\frac{1}{\beta-v}\Big{|}=\frac{\beta-\alpha}{|(\alpha-v)(\beta-v)|}$ (19)
see e.g. [6, p.10]. Taking $f(u)=\frac{t\log(1+\frac{u}{m})}{2\pi},a=0,b=m\varepsilon-1$ and $f^{\prime}(u)=\frac{t}{2\pi(m+u)}$ gives
$\sum_{-N\leq v\leq 0}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $\ll\sum_{-N\leq v\leq 0}\frac{\beta-\alpha}{(\alpha-v)(\beta-v)}$ (20)
$=(\beta-\alpha)\sum_{0\leq v\leq N}\frac{1}{(\alpha+v)(\beta+v)}$ (21)
$\ll(\beta-\alpha)\sum_{0\leq v\leq N}\frac{1}{1+v^{2}}$ (22)
$\ll\beta-\alpha$ (23)
$\sim\frac{t\varepsilon}{2\pi m}.$ (24)
Choose a rational $t\in[T,H]$ so that $\alpha=\frac{t}{2\pi(m+M)}$ and $\beta=\frac{t}{2\pi m}$ are never integers, where $m\in\Big{[}\frac{T}{(1+\varepsilon)\log^{2}T},T(1+\varepsilon)\Big{]}$ is an integer and $M$ is the largest integer $\leq m\varepsilon-1$. Thus $v\neq\alpha$ in particular. Let $K_{0}$ be the largest integer $<\alpha$. We also have
$\sum_{1\leq v\leq\alpha}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $=\sum_{1\leq v\leq K_{0}}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ (25)
$\ll\sum_{1\leq v\leq K_{0}}\frac{\beta-\alpha}{(\alpha-v)(\beta-v)}$ (26)
$\ll(\beta-\alpha)\int_{1}^{K_{0}}\frac{1}{(\alpha-v)(\beta-v)}\mathrm{d}v$ (27)
$\ll(\beta-\alpha)\log t$ (28)
$\ll(\beta-\alpha)\log t$ (29)
$\sim\frac{t\varepsilon}{2\pi m}\log t.$ (30)
Let $K_{1}$ be the least integer $>\beta$. Note that
$\sum_{\beta<v\leq N}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x$ $\ll\sum_{K_{1}\leq v\leq N}\frac{\beta-\alpha}{(v-\alpha)(v-\beta)}$ (31)
$\ll(\beta-\alpha)\int_{K_{1}}^{N}\frac{1}{(v-\alpha)(v-\beta)}\mathrm{d}v$ (32)
$\ll(\beta-\alpha)\log t$ (33)
$\ll(\beta-\alpha)\log t$ (34)
$\sim\frac{t\varepsilon}{2\pi m}\log t.$ (35)
Since $\beta-\alpha\sim\frac{t\varepsilon}{2\pi m}\ll\frac{\log^{2}T}{T^{1/4}}=o(1)$, note that the range $\alpha<v\leq\beta$ contains no integers. Thus
$\sum_{\alpha<v\leq\beta}\int_{a}^{b}e^{2\pi i(f(x)-vx)}\mathrm{d}x=0.$ (36)
Combining (36), (35), (30), (25) and (19), yields $\sum_{a<r\leq b}e^{2\pi if(r)}=o(1)$ and by Abel summation it follows that
$\sum_{a<r\leq b}\frac{1}{r}e^{2\pi if(r)}=\sum_{0<r\leq\varepsilon m-1}\frac{1}{r}e^{it\log(1+\frac{r}{m})}=o(1).$ (37)
Inserting (37) into the right-hand side of (18), summing over $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$ and applying the Prime Number Theorem that $\sum_{m\leq x}\Lambda(m)\sim x,$ yields
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (38)
For $\frac{m}{1+\varepsilon}\leq n\leq m-1$, put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ thus $\log(n/m)=-r/m(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-m/r(1+O(r/m))$. Thus by arguing as above, one arrives at
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{it}}{m^{1+\frac{1}{\log n}+it}\log(n/m)}=o(T).$ (39)
Summing (9) for $\frac{T}{\log^{2}T}<n<T$, changing the order of summation, invoking (39), (38), (14), (13) and (12) yields
$\sum_{\frac{T}{\log^{2}T}<n<T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+o(T)$ (40)
$=-\frac{T}{\pi}\log\log T+o(T)$ (41)
since $H=2T+O(1)$ and $\sum_{m\leq x}\frac{\Lambda(m)}{m}=\log x-B+o(1)$ where $B=0.57721\cdots$ denotes the Euler-Mascheroni constant [5, p.182]. By Lemma 2, we have
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{2\pi n}+O((\log T)\log\log 3n)$ (42)
uniformly for $2\leq n\ll T$. Define $N(y):=\sum_{0<\gamma\leq y}1$. Note that $N(y)=(y/2\pi)\log(y/2\pi)-y/2\pi+O(\log y)$ [5, Corollary 14.3]. Thus, summing (42) for $1\leq n\leq\frac{T}{\log^{2}T}$ gives
$\sum_{1\leq n\leq\frac{T}{\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(T)-\frac{T}{2\pi}\sum_{2\leq n\leq\frac{T}{\log^{2}T}}\frac{\Lambda(n)}{n}+o(T)$ (43)
$=\frac{T}{\pi}\log\log T+\frac{T}{2\pi}(B-1+\log 2)+o(T).$ (44)
Combining (44), (41), (8) and (3) yields
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n<T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (45)
$\gg T.$ (46)
Suppose that $\Theta<1$. Then by Lemma 1 we have
$\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta-1}$ (47)
uniformly for $T\leq\gamma_{T}\leq H.$ Hence
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}\ll T^{\Theta}\log T.$ (48)
But (48) is a contradiction to (46), thus our supposition must be false. This completes the proof. ∎
Corollary 1. There exists an infinite sequence of zeros of $\zeta$, whose real parts converge to 1.
Proof.
Combine Theorem 1 with the fact that $\zeta(s)\neq 0$ for $\Re(s)\geq 1$ [5, Theorem 6.6]. ∎
Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $T\leq\gamma_{T}\leq 2T+1$ and $\Re(\rho_{T})=1-o(1)$.
Remark 2. Our argument crucially relies on:
(i) The functional equation for $\zeta$, and
(ii) Theorem 4.11 of [7].
Both (i) and (ii) do not hold for Beurling zeta functions corresponding to Beurling number systems whose integers are not the ordinary integers. Hence our approach wouldn’t work for such, whose analogues of the RH are known to be true. Also, (ii) does not hold for the Weil zeta functions, which are known to satisfy the RH.
Remark 3.
From the literature on arXiv and elsewhere, it doesn’t appear like this approach on RH has been tried before. However, the idea of applying Gonek’s formula (42) in estimating sums over zeros of zeta in general, is not new. As one can confirm from the literature, it has proven to be a useful tool for such tasks. It is therefore instructive to sum (9) carefully, e.g. by not rushing to naively bound $n^{it}$ by 1 for $n$ not too close to $m$. Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$. Note that the analogue of (42) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{n^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite useful for our purposes.
Remark 4. The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need Theorem 4.11 of Titchmarsh [7] to also hold for $\sigma=\sigma_{t}=o(1)$ with an error term $\ll(\log T)^{-1}$. However, this is not possible (as can be seen from the approximate functional equation for $\zeta$). It can also be seen that Titchmarsh’s proof of his Theorem 4.11 doesn’t hold for $\sigma=\sigma_{t}=o(1)$. Indeed, he begins his argument by quoting the Euler-McLaurin summation formula for $\zeta(\sigma+it)$ (equation 3.5.3). This is initially defined for $\sigma>1$, and then is extended by holomorphic continuation to $\sigma>0$. Recall that a function can only be holomorphically extended on open sets, and that’s precisely why Titchmarsh assumes that $\sigma$ is bounded below by some constant $\sigma_{0}>0$.